Given a desired Future Worth accumulation, F = $16,000, an interest rate, i=5%, and a time period, n=7 years, calculate the unknown Present Worth of the single deposit, P=?

Please show all work and formulas.
Answer #1

Future value = 16000 $

Interest rate = 5%

Time or N = 7 Years

Present worth = Future value / (1 + R)^n

= 16000 / (1 + 5%)^7

= 16000 / (1 + 0.05)^7

= 16000 / (1.05)^7

= 16000 / 1.40710042265625

Present worth of deposit = 11370.90 or (11371 $ )